An object was thrown vertically upwards and it reached it peak height after 2.5 seconds before dropping back. What is the peak height measured from the throwing point? answer in meters
1
Expert's answer
2020-07-06T15:15:24-0400

The maximum height

"h_{max}=\frac{v_0^2}{2g}=\frac{gt^2}{2}""=\frac{9.8\:\rm m/s^2\times (2.5\: s)^2}{2}=30.6\:\rm m"